# Improving How Agents Cooperate: Attention Schemas in Artificial Neural Networks

### v1

@@ Introduction
A central hypothesis of the attention schema theory is that, in people, the presence of an attention schema allows for better social cognition and better performance in cooperative tasks [1,6]. In the present study, in Experiment 1, networks that contained attention schemas showed better ability to categorize each other’s attention states. We hypothesized that networks with attention schemas would also perform better on a joint task that requires interaction with another agent. We chose a simple task involving two networks mutually coloring an image, much like two children sharing a page in a coloring book. The networks took turns placing colored pixels on the image. Reward was earned jointly by both networks for painting as many pixels as possible over multiple turns, but reward was lost if the painted area of one network overlapped the painted area of the other network. One each turn, each agent chose which pixels to color, but did not have any direct information about which pixels its partner would choose on the same turn. Ideal performance on the task should require each network to make turn-by-turn predictions of the behavior of its partner. If one network could predict the part of the image that its partner was likely to color, it could then choose to color a different part of the image, avoiding overlap. Thus, mutual behavioral prediction should improve performance. The behavior of each network was in turn influenced by its visual attention to the underlying image. We therefore reasoned that it would be beneficial for each network to learn an implicit, predictive model of the other network’s attention.
@@
Figure 4A shows performance on the task across fifty training epochs, for all three types of network pairings. In line with our predictions, we found that schema-schema teams achieved the highest rewards on the cooperative coloring task (blue line in Figure 4A, averaged across epochs, mean = 2.04). Mixed teams received the second highest rewards (green line in Figure 4A, averaged across epochs, mean = 1.91). Control-control teams obtained the smallest rewards (orange line in Figure 4A, averaged across epochs, mean = 1.76).
Figure 4: Figure 4: Performance of network pairs on joint coloring task. A. Average reward accrued during the coloring of each image, as a function of training epoch. Shading shows 95% confidence interval. B. Average amount of overlapping pixels for each image, as a function of training epoch. Shading shows 95% confidence interval. C. Average number of pixels colored per image, for each network in the pair, as a function of training epoch. For the schema-schema pair, both networks had an attention schema. For the control-control pair, neither network had an attention schema. For the mixed pair, one network had an attention schema and the other did not. Shading shows 95% confidence interval.
We also analyzed the average number of overlapping pixels as a metric of within-team coordination, since networks that are well coordinated with their teammates should be better at reducing the overlap. The results are shown in Figure 4B. We found that schema-schema teams and mixed teams exhibited better coordination, selecting fewer overlapping pixels on average (blue line for schema-schema teams, mean across epochs = 54.51; green line for mixed teams, mean across epochs = 54.89), relative to the control-control teams (orange line for control control teams, mean across epochs = 57.79).
@@
The two green lines in Figure 4C show the result for the mixed, control-schema team. Here the behavior is markedly different. The solid green line shows the trajectory for the network that contained an attention schema. The average number of pixels selected by the network first dropped, then stayed low for the remainder of the epochs. This behavior suggests that the network was less well able to predict where its partner would place colored pixels. As a result, its best strategy was to learn to place relatively few pixels on each turn, to reduce the chance of blindly overlapping with its partner. In contrast, the dashed green line shows the trajectory for the network that did not contain an attention schema. The average number of pixels selected by the network rose steeply from epoch 10 to the end of training at epoch 50. This rise suggests that the network was successfully learning to predict where its partner would place colored pixels. As a result, its best strategy was to color increasingly larger numbers of pixels that it could correctly place in a manner that avoided overlapping with its partner. The data therefore suggest that in the mismatched pair, the schema network became more interpretable and predictable to the non-schema network.
@@
One of the properties of the joint coloring task that makes it useful in the present context is that on each turn, an agent has a large range of options. It chooses how many and which pixels to color. That choice is heavily shaped by the parts of the image that are enhanced by its attention mechanism. In this sense, the task resembles a typical human situation with open-ended choice. Should you reach for the coffee cup, the donut, the phone, or start a conversation? Moment by moment, the focus of your attention shifts to the feature of your world that will be the target of your next action. For this reason, being able to understand and predict attention is useful for understanding and predicting the behavior of others, and thus for cooperative interaction. In the joint coloring task, in a much simpler way, the same principles may be at work.
@@
It has been proposed that the human attention schema, a predictive model of attention, is important to our social cognition and our cooperative ability [1,6]. Using similar principles in artificial systems might be a way to improve the prosocial and cooperative behavior of machines. The goal of the present study was to test whether adding an attention schema to artificial agents would make them better cooperators in a joint task. We used a version of an attention schema in which a network learns to predict its own attention state and uses those predictions to modify its attention state [15].
In Experiment 1, we found that adding an attention schema to an agent improved its ability to learn how to categorize the attentional state of another agent. Even more so, it improved the ability to have its own attentional state categorized by others. This finding is consistent with recent work suggesting that when an agent learns a predictive self-model, it also partly self-regularizes [16]. It learns to make itself more predictable. One interpretation of the present results, therefore, is that an agent with an attention schema, by virtue of learning to predict its own attentional states, regularizes its own attentional dynamics and makes itself more easily categorized, interpreted, and predicted by others.
@@
In Experiment 3, we tested agents in a simple joint task, a coloring task. Two agents took turns coloring in an image while minimizing overlap. On each turn, each agent chose pixels to color without knowing the pixels chosen by its partner on that turn. Ideal performance should require each agent to predict the choices of its partner, such that it can maximize the number of pixels it colors while minimizing overlap with its partner’s choices. The earned rewards were shared by both agents. We found, as hypothesized, that performance on the task was best when both agents contained an attention schema, weaker when only one of the two agents contained an attention schema, and weakest when neither agent in the pair contained an attention schema. Further analysis suggested that the largest effect of adding an attention schema to an agent may have been to make that agent more predictable to its partner, allowing the partner to perform better at the task.
The study of social cognition has focused almost exclusively on how mechanisms in the brain of person A can allow that person to better understand person B. However, it may be equally important, if not more so, to consider the mechanisms in the brain of person A that render that person more easily understood by person B. It is a matter of not just rendering oneself a better receiver of socially useful information, but also a better sender of socially useful information. We speculate that possessing an attention schema may provide both properties – a greater ability to understand others, and a greater ability to be understood by others – thereby improving social cooperation. In a competitive interaction, it is probably disadvantageous to be more easily understood by others. But in a cooperative interaction, rendering oneself more transparent, more interpretable, and ultimately more predictable, may be an advantage.
The reason why an attention schema, in particular, may be of central importance, is that attention is almost entirely determinative of behavior. People react to items that they are paying attention to, and do not react to items they are not attending. Attention enhances the processing of an item and brings it to higher cognition and action planning. To ask what is in a person’s mind at any particular moment is essentially a colloquial way of asking what items are in the focus of that person’s attention. In a much simplified way, the present study might be seen as reflecting the same kind of processes.

### v2

@@ Introduction
A central hypothesis of the attention schema theory is that, in people, the presence of an attention schema allows for better social cognition and better performance in cooperative tasks [1,6]. The reason is that, at a simple, fundamental level, attention influences behavior, and therefore predicting attention state allows for better prediction of behavior. The purpose of the present set of studies was to examine how well the principles of attention schemas apply to a standard transformer-type artificial neural network. In experiment 1, networks that contained attention schemas showed better ability to categorize each other’s attention states. But are these networks actually better able to cooperate, if tested directly on a joint task? We hypothesized that networks with attention schemas would perform better on a joint task that requires interaction with another agent.
We chose a simple task involving two networks mutually coloring an image, much like two children sharing a page in a coloring book. The networks took turns placing colored pixels on the image. Reward was earned jointly by both networks for painting as many pixels as possible over multiple turns, but reward was decreased if the painted area of one network overlapped the painted area of the other network. On each turn, each agent chose which pixels to color. While each agent had information about its partner’s behavior on previous turns, it had no information about which pixels its partner would choose on the current turn. Ideal performance on the task should require each network to make turn-by-turn predictions of the behavior of its partner. If one network could predict the part of the image that its partner was likely to color, it could then choose to color a different part of the image, avoiding overlap. Thus, mutual behavioral prediction should improve performance. The behavior of each network was in turn influenced by its visual attention to the underlying image. We therefore reasoned that it would be beneficial for each network to learn an implicit, predictive model of the other network’s attention. Although a mutual coloring task, because of its simplicity, may not be an ideal test of social cooperation or mutual attention modeling in humans, it seemed, for the reasons just explained, to be well suited to testing the hypothesis in the artificial networks studied here.
@@
Figure 4A shows performance on the task across fifty training epochs, for all three types of network parings. In line with our predictions, we found that schema-schema teams achieved the highest rewards on the cooperative coloring task (blue line in Figure 4A, averaged across epochs, mean=2.04). Mixed teams received the second highest rewards (green line in Figure 4A, averaged across epochs, mean=1.91). Control-control teams obtained the smallest rewards (orange line in Figure 4A, averaged across epochs, mean=1.76).
Figure 4: Figure 4: Results for Experiment 3. Performance of network pairs on joint coloring task. A. Average reward accrued during the coloring of each image, as a function of training epoch. Shading shows 95% confidence interval. B. Average amount of overlapping pixels for each image, as a function of training epoch. Shading shows 95% confidence interval. C. Average number of pixels colored per image, for each network in the pair, as a function of training epoch. For the schema-schema pair, both networks had an attention schema. For the control-control pair, neither network had an attention schema. For the mixed pair, one network had an attention schema and the other did not. Shading shows 95% confidence interval.
We also analyzed the average number of overlapping pixels as a metric of within-team coordination, since networks that are well coordinated with their teammates should be better at reducing the overlap. The results are shown in Figure 4B. We found that schema-schema teams and mixed teams exhibited better coordination, selecting fewer overlapping pixels on average (blue line for schema-schema teams, mean across epochs=54.51; green line for mixed teams, mean across epochs=54.89), relative to the control-control teams (orange line for control-control teams, mean across epochs=57.79).
@@
The two green lines in Figure 4C show the result for the mixed, control-schema team. Here the behavior is markedly different. The solid green line shows the trajectory for the network that contained an attention schema. The average number of pixels selected by the network first dropped, then stayed low for the remainder of the epochs. This behavior suggests that of the two networks in the team, the network with an attention schema was less well able to predict where its partner would place colored pixels. As a result, its best strategy was to learn to place relatively few pixels on each turn, to reduce the chance of blindly overlapping with its partner. In contrast, the dashed green line shows the trajectory for the network that did not contain an attention schema. The average number of pixels selected by this network rose steeply from epoch 10 to the end of training at epoch 50. This rise suggests that the network was successfully learning to predict where its partner would place colored pixels. As a result, its best strategy was to color increasingly larger numbers of pixels which it could correctly place in a manner that avoided overlapping with its partner. The data therefore suggest that in the mismatched pair, the schema network became more interpretable and predictable to the non-schema network.
@@
It is important to reiterate that the artificial mechanism of attention used here, a transformer mechanism, is very different from the biological mechanism in the human brain, and that the task used here, though well suited to testing the artificial agent, might be too simple to provide an insightful test of human participants. However, there are shared underlying principles, and therefore we feel it is valid to speculate about how the present results may reflect on the human case. One of the properties of the joint coloring task that makes it useful in the present context is that on each turn, an agent has a large range of options. It chooses how many and which pixels to color. In that sense, the task loosely resembles a human situation with open-ended choice. Should you reach for the coffee cup, the donut, the phone, or start a conversation? Moment by moment, the focus of your attention shifts to the feature of your world that will be the target of your next action. For this reason, being able to understand and predict attention is useful for understanding and predicting the behavior of others, and thus modeling the attention of others should be useful for cooperative interaction. In the joint coloring task, in a much simpler way, the same principles appear to be at work. Future studies may investigate a more complex version of the coloring task in humans.
@@
It has been proposed that the human attention schema, a model of attention, is important to our social cognition and our cooperative ability [1,6]. The goal of the present study was to test whether adding an attention schema to artificial agents using a transformer-style of attention would make them better cooperators in a joint task. We tested the transformer architecture partly because, although the mechanism is quite different from the biological attention mechanism in the human brain, some of the underlying principles of feature selection and influence on behavior are similar; and partly because we hoped to test the effects of attention schemas on the most common and widely used attention mechanism in machine learning. We used a version of an attention schema in which a network learns to predict its own attention state and uses those predictions to modify its attention state [18].
In Experiment 1, we found that adding an attention schema to an agent improved its ability to learn how to categorize the attentional state of another agent. Even more so, it improved the ability to have its own attentional state categorized by others. This finding is consistent with recent work suggesting that when an agent learns a predictive self-model, it also partly self-regularizes [19]. It learns to make itself more predictable. One interpretation of the present results, therefore, is that an agent with an attention schema, by virtue of learning to predict its own attentional states, regularizes its own attentional dynamics and makes itself more easily categorized, interpreted, and predicted by others. We speculate that this insight is generalizable – that in biological systems as well, one of the main adaptive advantages of self-models may be to make the self more regularized, more predictable, and thus a better target for the social cognition of others. In that speculation, having a self-model makes one more predictable and interpretable to others, thereby enhancing social cooperativity.
@@
In Experiment 3, we tested agents in a simple joint task, a coloring task. Two agents took turns coloring in an image while minimizing overlap. On each turn, each agent chose pixels to color without knowing the pixels chosen by its partner on that turn. Ideal performance should require each agent to predict the choices of its partner, such that it can maximize the number of pixels it colors while minimizing overlap with its partner’s choices. The earned rewards were shared by both agents. We found, as hypothesized, that performance on the task was best when both agents contained an attention schema, weaker when only one of the two agents contained an attention schema, and weakest when neither agent in the pair contained an attention schema. Further analysis suggested that the largest effect of adding an attention schema to an agent may have been to make that agent more predictable to its partner, allowing the partner to perform better at the task. This may theoretically be achieved in any number of ways—making oneself more like one’s partner (thus, more predictable to the partner), making one’s moves simpler or less complicated, and so on.
The study of social cognition has focused almost exclusively on how mechanisms in the brain of person A can allow that person to better understand person B. However, it may be equally important, if not more so, to consider the mechanisms in the brain of person A that render that person more easily understood by person B. It is a matter of not just rendering oneself a better receiver of socially useful information, but also a better sender of socially useful information. We speculate that possessing an attention schema may provide both properties – a greater ability to understand others, and a greater ability to be understood by others – thereby improving social cooperation. In a competitive interaction, it is probably disadvantageous to be more easily understood by others. But in a cooperative interaction, rendering oneself more transparent, more interpretable, and ultimately more predictable, may be an advantage. We recognize that these suggestions emerge here out of a study of non-biological, artificial agents using artificial attention mechanisms, but we suggest that at least some of the underlying principles should be shared by all agents that use any form of attention, attention control, and attention schema.
The reason why an attention schema, in particular, may be of central importance, is that attention is almost entirely determinative of behavior. In the case of the artificial agents, attention determines the features that will most influence action. In the human case as well, people react to items that they are paying attention to, and do not react to items they are not attending. Human attention, the ability to control it, understand it, reveal it to others, and predict the attention of others, therefore plays a large role in social behavior. In a much-simplified way, the present study might be seen as reflecting the same kind of processes.